ABC Inc. is an all-equity firm with two million shares outstanding. The company is expected to have $16 million earnings and pay $6.4 million dividends next year. The other $9.6 million will be invested in projects with 20% return. It is expected that the company’s dividend payout ratio will remain the same for a foreseeable future. The company’s beta is estimated to be 1.2. Suppose the annualized yields on 91-day T-bills to be 5%. The analysts’ estimate of long-term risk premium for the S&P/TSX Comp. (market risk premium) is 15%.

a) What is the intrinsic (calculated) value of ABC Inc.’s share? Show your work.

b) How much of the value obtained in part (a) is attributed to the value of growth opportunities? Show your work.

c) If the market price of ABC is currently $60 and in one year, you expect the market price to be equal its intrinsic value at that time, what is the expected one-year holding period return (HPR) on ABC?
Answer #1

a). Expected earnings of Next year = $ 16 million

Expected Dividend to be paid next year = $ 6.4 million

Dividend Payout ratio = Expected Dividens/Expected earnings = $6.4 MM/$16 MM

= 40%

Since, Dividend payout ratio will remain same, calculating the growth rate

Growth rate(g) = ROE*plow back ratio

ROE = 20%

plow back ratio = 1- Dividend payout ratio = 1-0.40 = 0.60 or 60%

g= 20% * 60%

= 12%

Company's Beta = 1.2

Annualized yields on 91-day T-bills (risk free rate ) = 5%

Market risk premium = 15%

Calculating Required return on stock using CAPM,

Required return on stock = risk free rate + Beta(Market risk premium)

= 5% + 1.2 (15%)

= 23%

- Expected Dividend to be paid next year = $ 6.4 million

No of shares outstanding = 2 million

Dividend per share (D1) = $6.4 million/2 million = $3.2 per share

growth rate (g) = 12%

Required return on stock (ke) = 23%

Calculating intrinsic (calculated) value of ABC Inc.’s share using Dividend Growth model:

Intrinsic value = D1/(ke-g)

= $3.2(0.23-0.12)

= $ 29.09 per share

b). Value of share with no growth opportunity = D1/Ke

Intrinsic value of ABC's share with no growth opportunity = $ 3.2/0.23

= $ 13.91

Value attributable to growth opportunity = Intrinsic value - Intrinsic value with no growth opportunity

= $ 29.09 - $ 13.91

= $ 15.18 per share

c). Current market price of ABC = $ 60

Intrinsic Value price in One year = D2/(ke-g)

= 3.2(1+0.12)/(0.23-0.12)

= $ 32.58

Calculating the Expected one-year holding period return (HPR on ABC:

= [( Price in one year - Current price) + Current Dividend]/Current Price

= [($32.58-$60)+ $3.2]/$60

= -40.37%

So, expected one-year HPR is -40.37%

If you need any clarification regarding this solution, then you can ask in comments

If you like my answer then please Up-vote as it will be motivating